What is the initial volume of a 12.1mol/L hcl solution needed in order 500.0ml of a 2.5 mol/l solution of hcl


1
Expert's answer
2022-04-28T14:48:03-0400

The initial volume of the HCl 12.1 mol/L solution can be deduced from the statement of the moles number equality:

"n_1=n_2"

when "n" can be written in the following way:

"c_1V_1=c_2V_2"

"V_1=\frac{c_2V_2}{c_1}=\frac{2.5 mol/L \times 500.0 ml}{12.1 mol/L} = 103.3 ml"